OpenStudy (anonymous):

Simplify: (-2i+√5)(-2i-√-5)?

OpenStudy (cwrw238):

by distributive law (-2i+√5)(-2i-√-5) = -2i*-2i -2i*-(√-5) + √5* -2i + √5*(-√-5)

OpenStudy (cwrw238):

note i = √-1, so i^2 = -1

OpenStudy (cwrw238):

so -2i*-2i = 4 * i^2 = 4 * -1 = -4

OpenStudy (cwrw238):

and √-5 can be written √5 i

OpenStudy (anonymous):

so it be -4+ sqr(5) i

OpenStudy (anonymous):

no it just be plus 5..right?

OpenStudy (cwrw238):

no - i was just helping you do the tricky bits you must simplify what i wrote in my first post

OpenStudy (anonymous):

oh..